# Benchmarking LLM Faithfulness in RAG with Evolving Leaderboards

### v1

@@ 6.3 FaithJudge
Table 3 presents confusion matrices for both binary and ternary classification using FaithJudge. We observe that Benign summaries, in particular, are difficult for FaithJudge to classify correctly. In the ternary setting, FaithJudge often misclassifies Benign summaries, generally labelling them as Consistent. Similarly, Questionable summaries are classified unreliably, though this aligns with expectations. For this reason, we only employ FaithJudge for binary classification.
Finally, Figure 2 shows the sensitivity and specificity of FaithJudge as the number of annotated examples provided increases. Specificity remains consistently high, though slightly decreasing as more examples are given, while sensitivity notably improves as the number of examples increases. This indicates that providing more annotated examples causes FaithJudge to predict hallucinated cases more often and better identify hallucinations.
7 Leaderboard Rankings
Figure 3 compares the ranking of the 10 LLMs studied in FaithBench based on human-annotated hallucinations with rankings from FaithJudge and Vectara’s existing hallucination leaderboard.
The left-most plot shows that rankings vary depending on the type of hallucination considered: Unwanted, Benign, or Questionable, even when assessed by human annotation. The other plots show that when considering all types of hallucination annotations, rankings in FaithBench align more closely with FaithJudge than with the existing leaderboard. FaithJudge rankings show six inversions compared to rankings from FaithBench considering Unwanted, Benign, and Questionable hallucinations, while the existing leaderboard rankings using HHEM shows 16 inversions.
@@ 8 Conclusion
Beyond FaithBench, we extend FaithJudge to additional RAG tasks, including question answering and data-to-text generation, using annotated examples of hallucinations from the RAGTruth dataset. We discuss this further in Appendix A. We also apply FaithJudge to a broader set of LLMs, producing leaderboard-style rankings that currently include 30 models. We share some of these results in Appendix C, providing a framework for more accurate faithfulness evaluation across diverse models and RAG tasks. We hope to continue to update our leaderboard to evaluate new models and to use improved LLM judges.
@@ Acknowledgements
We respectfully acknowledge the late Simon Mark Hughes, who led the development of the original HHEM model and Vectara’s Hallucination Leaderboard. His contributions laid important groundwork for Vectara’s ongoing research and continue to leave a lasting influence on our work.
Limitations
There are some limitations with our evaluation methodology. First, our evaluation focuses exclusively on faithfulness and does not address the overall quality or usefulness of summaries and answers. Though summary and answer quality are important in RAG applications, we consider this evaluation somewhat orthogonal to faithfulness.
One issue to consider is that an extractive summarizer or an LLM that simply copies parts of or the entire article in its response would technically avoid hallucinations. Nonetheless, we maintain that evaluating LLMs through hallucinations in generated summaries is promising because these hallucinations remain persistent.
Finally, while the o3-mini-high judge demonstrates strong effectiveness, there remains room for enhancing accuracy and agreement with human annotators. We hope that as LLMs continue to improve, replacing o3-mini-high in FaithJudge would allow for more accurate and reliable evaluation.
Appendix A Adding More Evaluation Tasks
@@
While FaithBench provides hallucination annotations across 10 different LLMs, it is limited to evaluating summaries only.
To broaden the scope of FaithJudge beyond summarization, we incorporate annotated responses from the RAGTruth dataset Niu et al. (2024). RAGTruth includes three types of tasks: Summarization, Question Answering (QA), and a Data-to-Text generation task that requires generating an overview of a business from JSON data sourced from the Yelp Open Dataset. It contains human-annotated hallucination labels for responses generated by six LLMs: GPT-3.5, GPT-4 OpenAI (2023), Llama-2 (7B, 13B, 70B) Touvron et al. (2023), and Mistral-7B Jiang et al. (2023).
For each RAGTruth task, we take up to 150 sources (articles for summarization, queries and passages for question-answering, and JSON data for data-to-text) with their corresponding annotated responses from the test set first and then the dev set. We remove sources where none of the LLM responses have a hallucination annotation.
Table 4 compares the effectiveness of FaithJudge against the zero-shot FACTS Grounding JSON prompt previously shown as an effective prompt in Jacovi et al. (2025), on the FaithBench and RAGTruth subsets used in our leaderboard. In each setting, FaithJudge achieves stronger agreement with human hallucination annotations, highlighting its strength across tasks beyond summarization.
@@ Appendix B Judge Bias
The FACTS Grounding leaderboard Jacovi et al. (2025) uses three different LLM judges to mitigate bias arising from any single judge favoring its own outputs. Inspired by this, we analyze judge bias using Tables 5 and 6, which evaluate the impact of using different judges across all subsets included in our leaderboard.
Table 5 reports the effectiveness of three different LLMs when used as judges. The table shows that o3-mini-high remains a relatively effective LLM for FaithJudge, often scoring the highest. The table also shows that using multiple judges can improve effectiveness further, though we note that in some cases, individual LLMs can score higher than the majority vote approach between the three LLMs. For example, o3-mini-high scores higher than the ensembling approach when evaluating on the RAGTruth QA subset.
Table 6 explores how each judge model ranks other LLMs. Interestingly, o3-mini-high and llama-4-maverick both rank gemini-2.0-flash as having the fewest hallucinated responses, while gemini-2.0-flash ranks itself second to o3-mini-high, with only a small difference in counts (29 vs. 31).
While using multiple judges might enhance robustness and reduce individual model bias, we currently rely on a single judge to reduce computational costs. As stronger LLMs become available, we plan to update FaithJudge by substituting the current judge model with a more effective one.
Appendix C Leaderboard Rankings
Table 7 presents FaithJudge rankings for a range of LLMs. In addition to detecting hallucinations, we also prompt FaithJudge to flag responses that are invalid, for example, when a model fails to meaningfully summarize an article. For simplicity, we count these as hallucinated responses. Models are ranked based on their overall hallucination rate, calculated as the total number of hallucinated or invalid responses across all four evaluation subsets. We plan to continue evaluating LLMs using FaithJudge alongside the existing leaderboard.

### v2

@@ 6.3 Evaluating FaithJudge
Dataset FACTS Grounding Prompt FaithJudge Prompt
F1-Macro Balanced Accuracy F1-Macro Balanced Accuracy
RAGTruth-Data-to-text 77.1 75.1 86.3 85.1
RAGTruth-QA 76.9 81.6 83.4 85.4
RAGTruth-Summary 73.6 80.3 80.2 84.9
FaithBench-Summary 54.3 65.2 70.8 77.6
Table 4: Comparison between the FACTS Grounding zero-shot prompting approach (JSON Prompt) and the FaithJudge prompting approach on the subsets of data used in our leaderboard. In all cases we use an o3-mini-high LLM judge. For FaithJudge, we prompt the judge to evaluate LLM responses by providing the responses from the other LLMs in the dataset with their corresponding annotations. For FaithBench, we evaluate using all summaries, treating summaries labelled as Questionable or Benign as inconsistent summaries.
Table 3 presents confusion matrices for binary and ternary (including Benign labels) hallucination classification using FaithJudge. We observe that Benign summaries are difficult for FaithJudge to classify correctly. In the ternary setting, FaithJudge often misclassifies Benign summaries, generally labelling them as Consistent. Similarly, Questionable summaries are classified unreliably, though this aligns with expectations. For simplicity, we only employ FaithJudge for binary classification.
Figure 2 shows the sensitivity and specificity of FaithJudge as the number of annotated examples provided increases. Specificity remains consistently high, though slightly decreasing as more examples are given, while sensitivity notably improves as the number of examples increases. This indicates that providing more annotated examples leads FaithJudge to predict hallucinated cases more often and better identify hallucinations.
7 Adding More Evaluation Tasks
While FaithBench provides hallucination annotations across 10 different LLMs, it is limited to evaluating summaries only. To broaden the scope of FaithJudge beyond summarization, we incorporate annotated responses from the RAGTruth dataset ragtruth. RAGTruth includes three types of tasks: summarization, question-answering, and a data-to-text generation task that requires generating an overview of a business from JSON data sourced from the Yelp Open Dataset yelp_open_dataset_2021. RAGTruth contains human-annotated hallucination labels for responses generated by six different LLMs: GPT-3.5, GPT-4 gpt4, Llama-2 (7B, 13B, and 70B) touvron2023llama, and Mistral-7B jiang2023mistral7b.
For each RAGTruth task, we take up to 150 sources (articles for summarization, queries and passages for question-answering, and JSON data for data-to-text) with their corresponding annotated responses primarily from the test set and supplemented by the dev set where necessary. We remove sources where none of the LLM responses have a hallucination annotation.
Table 4 compares the effectiveness of FaithJudge against the zero-shot FACTS Grounding JSON prompt, which was previously shown to be an effective prompt in jacovi2025facts, on the FaithBench and RAGTruth subsets used in our leaderboard. In each setting, FaithJudge achieves stronger agreement with human hallucination annotations, highlighting its strength across tasks beyond summarization.
@@ 8 Conclusion
Our leaderboard is live on the FaithJudge GitHub, and we share some of these results in Appendix C, providing a framework for more accurate faithfulness evaluation across diverse models and RAG tasks. We plan to continue to update our leaderboard to evaluate new models and to use improved LLM judges.
Limitations
There are some limitations to our evaluation methodology. First, our evaluation focuses exclusively on faithfulness and does not address the overall quality or helpfulness of summaries and answers. Though summary and answer quality are important in RAG applications, we consider this evaluation largely orthogonal to faithfulness.
One issue to consider is that an extractive summarizer or an LLM that simply copies parts of the article or the entire article in its response would avoid hallucinations. Nonetheless, we maintain that evaluating LLMs through hallucinations in generated summaries is promising because these hallucinations remain persistent in current LLMs.
Finally, while the o3-mini-high judge demonstrates strong effectiveness, there remains room for enhancing accuracy and agreement with human annotators. We hope that as LLMs continue to improve, replacing o3-mini-high in FaithJudge may allow for more accurate and reliable evaluation.
@@ Acknowledgements
We respectfully acknowledge the late Simon Mark Hughes, who led the development of the original HHEM model and Vectara’s original hallucination leaderboard. His contributions laid important groundwork for Vectara’s ongoing research and continue to leave a lasting influence on our work.
Appendix
Appendix A Summaries with FaithJudge Verdicts
Source hong kong ( cnn ) six people were hurt after an explosion at a controversial chemical plant in china ’s southeastern fujian province sparked a huge fire , provincial authorities told state media. the plant , located in zhangzhou city , produces paraxylene ( px ) , a reportedly carcinogenic chemical used in the production of polyester films and fabrics  . the blast occurred at an oil storage facility monday night after an oil leak , though local media has not reported any toxic chemical spill . five out of six people were injured by broken glass and have been sent to the hospital for treatment , xinhua news agency reported . more than 600 firefighters were sent to battle the blaze and the fire is now under control , the zhangzhou fire department wrote on their official microblogging account tuesday morning . residents living close to the plant had heard the explosion and took to weibo to post photos of the fire . one user wrote that he heard a loud blast and felt slight tremors . the plant was hit by another explosion in july 2013 , although there were no reports of casualties or toxic leaks at the time . though demonstrations are illegal in china , the construction of px plants has sparked protests , which have occasionally turned violent , in several cities in recent years . the zhangzhou plant was slated for xiamen – a densely populated city in the southeast of the country . however , it provoked an angry backlash in 2007 due to pollution concerns and prompted the local government to relocate the factory to its current , more remote location .  Gemini-2.5-Pro Verdict: Inconsistent Summary An explosion at a chemical plant in Zhangzhou, China, injured six people and caused a large fire. The plant produces paraxylene (PX), a reportedly carcinogenic chemical. The blast occurred on a Monday night at an oil storage facility after an oil leak, though no toxic chemical spill has been reported. Five of the six people were injured by broken glass. Over 600 firefighters were sent to battle the blaze, which is now under control. The plant had a previous explosion in July 2013 and was relocated to its current location from Xiamen in 2007 due to public protests over pollution concerns. FaithJudge Response The summary accurately reflects almost all details from the article. It correctly states that the explosion injured six people, that five of them were injured by broken glass, that the blast occurred at an oil storage facility on Monday night after an oil leak, that no toxic chemical spill was reported, that over 600 firefighters controlled the fire, and that there was a previous explosion in July 2013. However, the summary’s last sentence introduces a key hallucination. Instead of stating that the plant “was slated for Xiamen” and that it “provoked an angry backlash in 2007 due to pollution concerns” (which then led the government to relocate the factory), the summary claims it “was relocated to its current location from Xiamen in 2007 due to public protests over pollution concerns.” This rephrasing mischaracterizes the source in two ways: 1. It implies the plant was actually located in Xiamen before being moved, rather than originally being planned there. 2. It changes “an angry backlash” to “public protests,” which is not how the source describes the reaction. Because these details are not supported by the source text, the summary contains hallucinations. Final classification: Inconsistent  Grok-3 Verdict: Inconsistent Summary An explosion at a chemical plant producing paraxylene (PX), a carcinogenic chemical used in polyester production, in Zhangzhou city, Fujian province, China, injured six people, five of whom were hurt by broken glass and hospitalized. The blast, caused by an oil leak at an oil storage facility, occurred on Monday night and led to a massive fire, which was controlled by over 600 firefighters by Tuesday morning. No toxic chemical spill was reported. Local residents heard the explosion, felt tremors, and shared images of the fire on Weibo. This plant, previously the site of a 2013 explosion with no casualties or leaks, was relocated from Xiamen to its current remote location due to pollution concerns and public backlash in 2007. PX plants have been controversial in China, sparking protests, sometimes violent, despite demonstrations being illegal. FaithJudge Response This summary introduces several details that do not exactly match the source. For example, it describes PX as “a carcinogenic chemical” instead of “a reportedly carcinogenic chemical,” changing the certainty level. It also states that the plant “was relocated from Xiamen” rather than noting that it was originally slated for Xiamen before the backlash prompted relocation. In addition, the summary phrases the 2013 explosion as having “no casualties or leaks” rather than saying there were “no reports” of casualties or toxic leaks. These deviations, even if minor, represent hallucinated details. Final classification: Inconsistent
Figure 3: Two LLM summaries of a news source from FaithBench with FaithJudge hallucination judgements. The red text indicates portions of the summaries that contain hallucinations.
Figure 3 shows examples of two different LLM summaries for a news source from FaithBench. We generally observe that hallucinations can often be subtle and require great effort to identify. For example, while the source identifies Paraxylene as a “reportedly carcinogenic chemical”, Grok-3 drops the “reportedly” modifier and describes Paraxylene as a “carcinogenic chemical”. This removed detail is pertinent because it changes the information from the source. Notably, to the best of our understanding, there is not enough evidence to say that Paraxylene is carcinogenic in humans or in animals kandyala2010xylene.
Dataset Judge Model F1-Macro Balanced Accuracy
RAGTruth (Data-to-text) o3-mini-high 86.3 85.1
gemini-2.0-flash 83.6 84.0
llama-4-maverick 82.1 80.6
Majority Vote 86.4 85.8
RAGTruth (QA) o3-mini-high 83.4 85.4
gemini-2.0-flash 81.8 84.2
llama-4-maverick 77.5 81.2
Majority Vote 81.0 83.8
RAGTruth (Summary) o3-mini-high 80.2 84.9
gemini-2.0-flash 83.6 82.7
llama-4-maverick 78.0 83.7
Majority Vote 84.6 88.0
FaithBench (Summary) o3-mini-high 70.8 77.6
gemini-2.0-flash 66.1 75.5
llama-4-maverick 74.7 76.9
Majority Vote 72.4 79.1
Table 5: Evaluation results for three models and an ensemble approach on the subsets of data used in our leaderboard. Here, for FaithBench, we evaluate using all summaries, treating summaries labelled as Questionable or Benign as inconsistent summaries.
@@ Appendix B Judge Bias
The FACTS Grounding leaderboard jacovi2025facts uses three different LLM judges to mitigate bias arising from any single judge favoring its own outputs. Inspired by this, we analyze judge bias using Tables 5 and 6, which evaluate the impact of using different judges across all subsets included in our leaderboard.
Table 5 reports the effectiveness of three different LLMs when used as judges. The table shows that o3-mini-high remains a relatively effective LLM for FaithJudge, often scoring the highest. The table also shows that while using multiple judges can improve effectiveness further, in some cases, individual LLMs can score higher than the majority vote approach between the three LLMs. For example, o3-mini-high scores higher than the ensembling approach when evaluating on the RAGTruth QA subset.
Table 6 explores how each judge model ranks the other LLMs. Interestingly, o3-mini-high and llama-4-maverick both rank gemini-2.0-flash as having the fewest hallucinated responses, while gemini-2.0-flash ranks itself second to o3-mini-high, with only a small difference in hallucinated response counts (29 vs. 31).
While using multiple judges might enhance robustness and reduce individual model bias, we currently rely on a single judge to reduce computational costs. As newer and stronger LLMs become available, we plan to update FaithJudge by substituting the current o3-mini-high judge model with a more effective one.
Appendix C FaithJudge Rankings
Table 7 presents FaithJudge rankings for a small set of LLMs. In addition to detecting hallucinations, we also prompt FaithJudge to flag responses that are invalid, for example, when a model fails to meaningfully summarize an article. For simplicity, since these cases are rare, we count these as hallucinated responses. Models are ranked based on their overall hallucination rate, calculated as the total number of hallucinated or invalid responses across all four evaluation subsets. We plan to continue evaluating LLMs using FaithJudge. Please see the current leaderboard, with many more LLMs evaluated, at https://github.com/vectara/FaithJudge.